OpenStudy (anonymous):

How do you find the centres and radii of circles with the base equation of ax^2 + ay^2 + bx + cy+ d=0???

OpenStudy (campbell_st):

complete the square in x and y \[ax^2 + bx + (b/2)^2 + y^2 +cy + (c/2)^2 = d + (b/2)^2 + (c/2)^2\]

OpenStudy (shamim):

wanna give u an attachment

OpenStudy (shamim):

because i m frm android cell phone

OpenStudy (shamim):

difficult to type equation

OpenStudy (anonymous):

ok

OpenStudy (shamim):

u know equation of a circle is (x-h)^2+(y-k)^2=a^2

OpenStudy (shamim):

h,k=coordinates of center

OpenStudy (shamim):

a=radius of circle

OpenStudy (shamim):

r u able to devide ur given equation by a

OpenStudy (shamim):

write wt u will get aftet devide by a

OpenStudy (amistre64):

i have an android from metropcs ... this site doesnt work worth a hill of beans on it